# SAMA-UNet: Enhancing Medical Image Segmentation with Self-Adaptive Mamba-Like Attention and Causal-Resonance Learning

### v1

@@ Abstract
Medical image segmentation plays an important role in various clinical applications, but existing models often struggle with the computational inefficiencies and challenges posed by complex medical data. State Space Sequence Models (SSMs) have demonstrated promise in modeling long-range dependencies with linear computational complexity, yet their application in medical image segmentation remains hindered by incompatibilities with image tokens and autoregressive assumptions. Moreover, it is difficult to achieve a balance in capturing both local fine-grained information and global semantic dependencies. To address these challenges, we introduce SAMA-UNet, a novel architecture for medical image segmentation. A key innovation is the Self-Adaptive Mamba-like Aggregated Attention (SAMA) block, which integrates contextual self-attention with dynamic weight modulation to prioritise the most relevant features based on local and global contexts. This approach reduces computational complexity and improves the representation of complex image features across multiple scales. We also suggest the Causal-Resonance Multi-Scale Module (CR-MSM), which enhances the flow of information between the encoder and decoder by using causal resonance learning. This mechanism allows the model to automatically adjust feature resolution and causal dependencies across scales, leading to better semantic alignment between the low-level and high-level features in U-shaped architectures. Experiments on MRI, CT, and endoscopy images show that SAMA-UNet performs better in segmentation accuracy than current methods using CNN, Transformer, and Mamba. The implementation is publicly available at GitHub.
@@
journal: Journal
[…]
[label1]organization=Faculty of Computing and IT, Sohar University, city=Sohar, postcode=311, country=Oman
[…]
[label2]organization=Division of Robotics, Perception and Learning(RPL), Department of Intelligent Systems, addressline=KTH Royal Institute of Technology, city=Stockholm, postcode=10044, country=Sweden
[…]
@@ 1 Introduction
Transformers, originally developed for natural language processing (NLP), have been successfully adapted to computer vision tasks, including medical image segmentation. Vision Transformers (ViT) [3] and Swin Transformers [4] are excellent at capturing global dependencies because they interpret images as sequences of tokens. Hybrid models such as TransUNet [5] and SwinUNETR [6] combine the local feature extraction of CNNs with the global modeling strengths of Transformers that lead to improved segmentation performance. However, Transformers produce computational complexity with respect to the sequence length, which becomes prohibitive when dealing with high-resolution medical images that require processing long sequences. State Space Sequence Models (SSMs) [7, 8] have gained significant attention for their ability to model long-range dependencies with linear computational complexity. These models have achieved success in NLP tasks and have been extended to computer vision, where approaches like Visual State Space (VSS) [9], bi-directional SSM [10], and omnidirectional SSM [11] have demonstrated promising results. Studies have highlighted that the Mamba module [12], a macro design within SSMs, significantly contributes to feature extraction in image-based tasks. However, while Mamba has achieved success in medical image segmentation, it also faces inherent difficulties. Specifically, Yu and Wang [13] has presented Mamba’s inability to incorporate causal relationships between image tokens limits its application in autoregressive tasks such as segmentation.
@@
To address these limitations, we introduce SAMA-UNet, a novel architecture designed to integrate Mamba-inspired innovations into transformer-based models. SAMA-UNet presents the SAMA block that combines contextual self-attention with dynamic feature modulation to enhance both local and global feature representations. This approach improves segmentation accuracy and reduces the computational complexity traditionally associated with transformers. Furthermore, we propose the CR-MSM, which enhances the integration of multi-scale features by introducing causal resonance learning. This module leverages the inherent multi-scale feature generation in U-shaped networks, preserving causal dependencies while optimizing feature representation across scales.
We aim to address the quadratic complexity of transformers by incorporating softmax attention mechanisms inspired by the human foveal vision system to enable better attention distribution across both localized and global contexts. Our goal is further enhanced by differential attention, which suppresses irrelevant noise tokens, and flash attention, which improves the efficiency of the attention mechanism. These innovations allow SAMA-UNet to achieve high-performance segmentation while maintaining computational efficiency. In this paper, we propose the following contributions:
  • 1.
    We present the SAMA Block that integrates dynamic feature modulation and attention mechanisms to achieve linear complexity while capturing both local and global feature dependencies.
  • 2.
    We suggest the CR-MSM to enhance feature fusion in U-shaped networks and improve the consistency of multi-scale information flow.
  • 3.
@@ 2.1 Transformers-based methods
Transformers have become a leading architecture in natural language processing (NLP), computer vision, and multimodal tasks due to their ability to capture long-range dependencies and represent complex features. However, the quadratic computational complexity of traditional transformer models when processing high-resolution images remains a significant challenge. Researchers have proposed several methods to mitigate this issue while maintaining performance [14]. Sparse attention techniques aim to reduce computational complexity by shortening token sequences. For example, shifted window-based attention [4, 15] computes attention maps within smaller windows and extends global attention coverage through overlapping sliding windows. Similarly, Dilated Attention [16] employs dilated rates to capture short- and long-range dependencies with limited sampling points. Deformable attention [17, 18, 19] dynamically adjusts sampling positions based on learnt offsets to allow more flexible attention allocation tailored to specific input features. Furthermore, pixel-focused attention simulates how human vision works, allowing each feature point to see both local and global details. In addition, linear attention replaces the softmax operation in standard attention with a non-negative mapping function to reduce the complexity from quadratic to linear time. However, such an approach often leads to performance degradation. Building on this, focused linear attention [20] improves the representations of the attention matrix by designing efficient mapping functions. Additionally, mamba-like linear attention [12] combines SSM and linear attention in a larger design inspired by Mamba, which further boosts performance. Finally, flash attention [21] makes the data transfer and computation processes for softmax attention on GPUs faster and more efficient, while still maintaining good performance.
@@
These advancements have inspired us to develop SAMA-UNet, which combines Mamba-like macro-architectural design into pixel-focused attention by using hardware-accelerated differential attention. This presented approach enables self-attention to achieve more efficient computational complexity while preserving high representation capacity, which is especially for high-resolution medical image segmentation tasks.
@@ 2.2 Mamba-based methods
State Space Sequence Models [7, 8] have shown promising results in modelling long-range dependencies with linear computational complexity in NLP and computer vision. The Mamba module, a specific SSM-based method, introduces a selective mechanism that enhances context-based reasoning by enabling input-dependent interactions along sequences via linear transformations. Mamba’s macro-architectural design has been applied in hybrid CNN-SSM structures to improve image processing tasks, particularly in medical image segmentation [22]. Several extensions of Mamba have been proposed to improve segmentation performance. For example, U-Mamba [23] integrates Mamba blocks after convolutional layers into the U-Net encoder, outperforming traditional CNN and transformer-based models. Swin-UMamba [24] and VM-UNet [25] replace convolutional layers in U-Net with components inspired by Visual State Space, improving the representation of local and global features. VM-UNet-V2 [26] refines skip connections using the convolutional block attention mechanism and introduces a Semantics and Detail Infusion module to strengthen the interaction between low- and high-level features. At the same time, LKM-UNet [27] uses patch-based methods along with Mamba modules to extract features at both local and global levels. Additionally, MSVM-UNet [28] adds multi-scale convolutions into Visual State Space blocks to effectively capture 2D features at different scales.
@@ 3.3 CR-MSM
The CR-MSM is designed to address the challenges of multi-scale feature integration, especially in U-shaped architectures. Medical images often exhibit multi-scale information that must be effectively captured to ensure accurate segmentation. CR-MSM introduces causal resonance learning, which ensures that the model maintains causal dependencies across different spatial scales while optimizing the flow of information between the encoder and decoder.
Unlike traditional multi-scale modules that rely solely on convolutions, CR-MSM explicitly captures directionality using multi-view transformations and State Space Modeling (SSM). In this context, SSM refers to a class of efficient sequence modeling techniques that treat spatial feature maps as 1D sequences and use structured linear recurrence to capture long-range dependencies. SSM models make them suitable for modeling spatial sequences over flattened and transformed feature maps, without the compute burden of attention mechanisms.
Given a set of feature maps from the encoder level, CR-MSM applies three spatial transformations—original, transposed, and flipped—to generate diverse directional views. These are flattened and concatenated across scales to create three distinct sequence channels. Each directional sequence is passed through a SSM block, capturing long-range contextual relationships across spatial orientations. The resulting sequences are then split, inverse-transformed, and reconstructed to their original format. Finally, we average the three directional results and send them through a linear projection layer, resulting in a strong and well-aligned multi-scale representation. We pass the output of CR-MSM to the decoder for fusion, which helps bridge the semantic gap between the encoder and decoder, thereby improving resolution-aware semantic consistency.
@@ 4.1.2 Evaluation metrics
Here, $S_{P}$ denotes the set of surface points from the predicted segmentation, and $S_{G}$ denotes the set of surface points from the ground truth segmentation. The function $\text{dist}(x,S)$ represents the shortest distance from point $x$ to the surface $S$. The parameter $\tau$ is a predefined tolerance value, and ||$\cdot$|| ⋅ | denotes the cardinality, i.e., the number of points in the set.
@@ 4.2 Quantitative and Qualitative results
On the BTCV dataset, our model outperforms the second-best model U-Mamba by 0.92% and 1.15% in DSC and NSD, respectively. We follow U-Mamba’s evaluation standard, eight out of thirteen organ classes (Aorta, Gallbladder, Kidney (L), Kidney (R), Liver, Pancreas, Spleen, and Stomach) are included in the calculation. Mamba based methods demonstrate strong overall performance, with our model maintaining a lead of 0.92% and 1.15% over U-Mamba in DSC and NSD. For the ACDC dataset, where all methods achieve relatively high accuracy, our method still surpasses SwineUNETR by 0.32% and 0.20% in DSC and NSD, respectively. On the EndoVis17 dataset, where both CNN-based and Mamba-based methods perform well, our method achieves further improvements, surpassing U-Mamba by 1.69% in DSC and 1.84% in NSD. Finally, on the ATLAS23 dataset, our model secures improvement by 1.27% and 1.68% in DSC and NSD, respectively while compare with U-Mamba.

### v2

@@ Abstract
Medical image segmentation plays an important role in various clinical applications, but existing deep learning models face trade-offs between efficiency and accuracy. Convolutional Neural Networks (CNNs) capture local details well but miss global context, whereas Transformers handle global context but at high computational cost. Recently, State Space Sequence Models (SSMs) have recently shown potential for capturing long-range dependencies with linear complexity, but their direct use in medical image segmentation remains limited due to incompatibility with image structures and autoregressive assumptions. To overcome these challenges, we propose SAMA-UNet, a novel U-shaped architecture that introduces two key innovations. First, the Self-Adaptive Mamba-like Aggregated Attention (SAMA) block adaptively integrates local and global features through dynamic attention weighting, enabling efficient representation of complex anatomical patterns. Second, the Causal-Resonance Multi-Scale Module (CR-MSM) improves encoder–decoder interactions by adjusting feature resolution and causal dependencies across scales, enhancing semantic alignment between low- and high-level features. Extensive experiments on MRI, CT, and endoscopy datasets demonstrate that SAMA-UNet consistently outperforms CNN-, Transformer-, and Mamba-based methods. For example, it achieves 85.38% DSC and 87.82% NSD on BTCV, 92.16% and 96.54% on ACDC, 67.14% and 68.70% on EndoVis17, and 84.06% and 88.47% on ATLAS23, establishing new benchmarks across modalities. These results confirm the effectiveness of SAMA-UNet in combining efficiency with accuracy, making it a promising solution for real-world clinical segmentation tasks. The source code is available at GitHub.
@@ 1 Introduction
Transformers, originally developed for natural language processing (NLP), have been successfully adapted to computer vision tasks, including medical image segmentation. Vision Transformers (ViT) [3] and Swin Transformers [4] are excellent at capturing global dependencies because they interpret images as sequences of tokens. Hybrid models such as TransUNet [5] and SwinUNETR [6] combine the local feature extraction of CNNs with the global modeling strengths of Transformers that lead to improved segmentation performance. However, Transformers produce computational complexity with respect to the sequence length, which becomes prohibitive when dealing with high-resolution medical images that require processing long sequences. State Space Sequence Models (SSMs) [7, 8] have gained significant attention for their ability to model long-range dependencies with linear computational complexity. These models have achieved success in NLP tasks and have been extended to computer vision, where approaches like Visual State Space (VSS) [9], bi-directional SSM [10], and omnidirectional SSM [11] have demonstrated promising results. Studies have highlighted that the Mamba module [12], a macro design within SSMs, significantly contributes to feature extraction in image-based tasks. However, while Mamba has achieved success in medical image segmentation, it also faces inherent difficulties. Specifically, Yu and Wang [13] has presented Mamba’s inability to incorporate causal relationships between image tokens limits its application in autoregressive tasks such as segmentation. Yu and Wang [13] highlighted that one of the key limitations of Mamba is its inability to incorporate causal relationships between image tokens, which restricts its applicability to autoregressive tasks such as segmentation. In medical image segmentation the causal relationship can be understood in a spatial and hierarchical way where fine details at higher resolutions naturally come first and support the formation of coarser semantic features. Preserving this order is important to keep encoder and decoder features consistent. Recent studies have emphasized the importance of advanced feature fusion strategies to improve segmentation accuracy in complex medical imaging scenarios. Ju et al. [14] proposed CLDSINet, a collaborative learning framework that jointly models static and dynamic information using multi-branch feature extraction to enhance feature representation. This development motivates our design of the CR-MSM module, which explicitly preserves hierarchical dependencies across scales while maintaining computational efficiency. Similarly, Multi-View Convolutional Networks (MVCNs) [15] have demonstrated the benefit of combining multiple directional or perspective views for improved feature learning, but their reliance on parallel CNN branches makes them computationally expensive for high-resolution segmentation tasks.
@@
To address these limitations, we introduce SAMA-UNet, a novel architecture designed to integrate Mamba-inspired innovations into transformer-based models. SAMA-UNet presents the SAMA block that combines contextual self-attention with dynamic feature modulation to enhance both local and global feature representations. This approach improves segmentation accuracy and reduces the computational complexity traditionally associated with transformers. Furthermore, we propose the CR-MSM, which enhances the integration of multi-scale features by introducing causal resonance learning where fine-grained encoder features resonate with semantically richer decoder features through a controlled multi-scale fusion process. This ensures that hierarchical dependencies across scales are not disrupted, thereby improving feature alignment and overall segmentation accuracy. This module encourages the inherent multi-scale feature generation in U-shaped networks to preserve causal dependencies while optimizing feature representation across scales. In addition, we address the quadratic complexity of transformers by incorporating a softmax attention mechanism inspired by the human foveal vision system. In human vision, the fovea prioritizes high-resolution detail in the center of focus while processing peripheral information more broadly and efficiently [16]. Inspired by this biological principle, our pixel-focused attention dynamically allocates attention weights to give more importance to local fine details while still capturing global context and reducing unnecessary computations. We further employ differential attention to suppress irrelevant noise tokens and flash attention to accelerate efficiency on modern hardware. These innovations collectively allow SAMA-UNet to achieve high-performance segmentation while maintaining computational efficiency. In this paper, we propose the following contributions:
  • We present the SAMA Block, inspired by human foveal vision. Splitting channels into local-neighborhood and condensed global context branches enables linear computational complexity, while differential attention suppresses irrelevant tokens. By integrating Mamba’s macro-structure, SAMA maximizes channel-wise information use to yield richer feature representations.
  • We suggest the CR-MSM to enhance feature fusion in U-shaped networks and improve the consistency of multi-scale information flow. By adjusting the SS2D token-scanning strategy to preserve causal order across scales, CR-MSM aligns with the autoregressive assumption. This design facilitates the complementary information among the various scales and effectively narrows the semantic gap between encoder and decoder features.
@@ 2.1 Transformers-based methods
Transformers have become a leading architecture in natural language processing (NLP), computer vision, and multimodal tasks due to their ability to capture long-range dependencies and represent complex features. However, the quadratic computational complexity of traditional transformer models when processing high-resolution images remains a significant challenge. Researchers have proposed several methods to mitigate this issue while maintaining performance [17]. Sparse attention techniques aim to reduce computational complexity by shortening token sequences. For example, shifted window-based attention [4, 18] computes attention maps within smaller windows and extends global attention coverage through overlapping sliding windows. Similarly, Dilated Attention [19] employs dilated rates to capture short- and long-range dependencies with limited sampling points. Deformable attention [20, 21, 22] dynamically adjusts sampling positions based on learnt offsets to allow more flexible attention allocation tailored to specific input features. Furthermore, pixel-focused attention simulates how human vision works, allowing each feature point to see both local and global details. In addition, linear attention replaces the softmax operation in standard attention with a non-negative mapping function to reduce the complexity from quadratic to linear time. However, such an approach often leads to performance degradation. Building on this, focused linear attention [23] improves the representations of the attention matrix by designing efficient mapping functions. Additionally, mamba-like linear attention [12] combines SSM and linear attention in a larger design inspired by Mamba, which further boosts performance. Finally, flash attention [24] makes the data transfer and computation processes for softmax attention on GPUs faster and more efficient, while still maintaining good performance. Flash attention [24] accelerates GPU computation for softmax attention, improving efficiency while maintaining accuracy.
Transformers have become a leading architecture in natural language processing (NLP), computer vision, and multimodal tasks due to their ability to capture long-range dependencies and represent complex features. However, the quadratic computational complexity of traditional transformer models when processing high-resolution images remains a significant challenge. Researchers have proposed several methods to mitigate this issue while maintaining performance [17]. Sparse attention techniques aim to reduce computational complexity by shortening token sequences. For example, shifted window-based attention [4, 18] computes attention maps within smaller windows and extends global attention coverage through overlapping sliding windows. Similarly, Dilated Attention [19] employs dilated rates to capture short- and long-range dependencies with limited sampling points. Deformable attention [20, 21, 22] dynamically adjusts sampling positions based on learnt offsets to allow more flexible attention allocation tailored to specific input features. Pixel-focused attention simulates how human vision works, allowing each feature point to see both local and global details. In addition, linear attention replaces the softmax operation in standard attention with a non-negative mapping function to reduce the complexity from quadratic to linear time, although often at the expense of performance. Building on this, focused linear attention [23] improves the representation of the attention matrix by designing efficient mapping functions. Mamba-like linear attention [12] combines SSM and linear attention in a larger design inspired by Mamba, further boosting performance. Flash attention [24] accelerates GPU computation for softmax attention, improving efficiency while maintaining accuracy. In addition to these attention optimizations, recent works have also emphasized adaptive feature fusion to improve efficiency. For instance, AFBNet [25] proposed a lightweight adaptive feature fusion module that balances local and global context, while DeepU-Net [26] employed a dual-branch design for multiscale fusion in remote sensing tasks. Such approaches highlight the trend toward efficient integration of features across scales, which motivates the design of our CR-MSM module.
@@
These advancements have inspired us to develop SAMA-UNet, which combines Mamba-like macro-architectural design with pixel-focused attention and hardware-accelerated differential attention. This enables self-attention to achieve more efficient computational complexity while preserving high representation capacity, particularly for high-resolution medical image segmentation.
2.2 Hybrid CNN–Transformer methods
CNNs have been the backbone of medical image segmentation because of their strength in capturing local spatial features. Classical models such as U-Net [2] and its variants U-Net++ [27], Attention U-Net [28], and Dense U-Net [29] achieve strong performance through hierarchical feature extraction and skip connections that improve fine-grained localization. However, CNNs are limited by their local receptive fields and struggle to capture long-range dependencies in high-resolution images.
To address this, hybrid CNN–Transformer models have emerged, combining CNNs’ local detail extraction with Transformers’ global context modeling. TransUNet [5] embedded a Transformer encoder into U-Net, while SwinUNETR [6] used hierarchical Swin Transformer blocks for multi-scale representation. More recently, methods such as EFF-ResNet-ViT [30], ResLNet[31], and DCSSGA-UNet [32] have explored efficient CNN–Transformer hybrids that integrate attention and fusion mechanisms to selectively enhance critical features while reducing redundancy. These hybrids achieve a balance between local and global learning but often introduce high computational cost. Our proposed SAMA-UNet builds on this line of work by embedding a self-adaptive Mamba-like attention mechanism into a CNN–Transformer-inspired design. Unlike existing hybrids, SAMA-UNet explicitly models causal dependencies across scales with the CR-MSM module, enabling more efficient feature alignment while maintaining computational efficiency.
@@ 2.3 Mamba-based methods
State Space Sequence Models [7, 8] have shown promising results in modelling long-range dependencies with linear computational complexity in NLP and computer vision. The Mamba module, a specific SSM-based method, introduces a selective mechanism that enhances context-based reasoning by enabling input-dependent interactions along sequences via linear transformations. Mamba’s macro-architectural design has been applied in hybrid CNN-SSM structures to improve image processing tasks, particularly in medical image segmentation [33]. Several extensions of Mamba have been proposed to improve segmentation performance. For example, U-Mamba [34] integrates Mamba blocks after convolutional layers into the U-Net encoder, outperforming traditional CNN and transformer-based models. Swin-UMamba [35] and VM-UNet [36] replace convolutional layers in U-Net with components inspired by Visual State Space, improving the representation of local and global features. VM-UNet-V2 [37] refines skip connections using the convolutional block attention mechanism and introduces a Semantics and Detail Infusion module to strengthen the interaction between low- and high-level features. At the same time, LKM-UNet [38] uses patch-based methods along with Mamba modules to extract features at both local and global levels. Other recent CNN–attention hybrid designs have also targeted the semantic gap problem. For example, DCSSGA-UNet [32] combines DenseNet201 with channel–spatial attention (CSA) and semantic guidance attention (SGA) to selectively enhance critical features while reducing redundancy between encoder and decoder pathways. Additionally, MSVM-UNet [39] and MSM-UNet [40] adds multi-scale convolutions into Visual State Space blocks to effectively capture 2D features at different scales.
@@ 3.3 CR-MSM
The CR-MSM is designed to enhance multi-scale feature fusion in SAMA-UNet by explicitly preserving the natural order of information flow across scales. Traditional fusion strategies, such as simple concatenation and convolutional aggregation, often mix high- and low-level features without regard to their hierarchical relationship, which disrupts semantic consistency between the encoder and decoder. CR-MSM addresses this limitation through causal resonance learning, where information from fine-grained features resonates with semantically richer features at coarser scales, ensuring that feature fusion reinforces rather than distorts the natural progression from detail to abstraction. Unlike MVCNs, which require parallel CNN branches or external multi-view inputs to capture diverse perspectives, CR-MSM generates directional views internally from the same feature map and fuses them efficiently using State Space Modeling. This design avoids the heavy computational overhead of MVCNs while achieving more consistent feature alignment.
In U-shaped networks such as UNet, encoder features at higher resolutions naturally precede and support the formation of lower-resolution features. This hierarchical dependency forms a causal chain where local fine details act as the cause and global semantic features are the effect. CR-MSM explicitly models this causal dependency by aligning encoder and decoder features so that the structural flow of information is preserved. By doing so, the module narrows the semantic gap across scales and improves segmentation accuracy, especially for complex anatomical patterns. To achieve this, CR-MSM introduces directional multi-view transformations combined with SSM. Each encoder feature map is transformed in four orientations which are original, transposed, flipped, and flipped-transposed to capture spatial dependencies from multiple perspectives. These directional views are then flattened in 1D sequences and individually processed through an SSM block, which employs structured linear recurrence to model long-range dependencies with linear computational complexity. This design allows CR-MSM to efficiently capture contextual relationships across orientations and avoiding the quadratic cost of self-attention while retaining strong representational power.
After SSM processing, the outputs of the four directional views are reconstructed to their original orientation and fused through causal resonance averaging. Unlike simple concatenation, which may merge features indiscriminately, causal resonance averaging respects the inherent order of feature generation and ensures that fine details and abstract features reinforce one another. Formally, for feature maps $F_{1},F_{2},\ldots,F_{N}$ at different scales, the transformed sequences $X_{1},X_{2},X_{3},X_{4}$ are processed into outputs $Y_{1},Y_{2},Y_{3},Y_{4}$ by the SSM, and the final fused feature map is given by:
$Y_{\text{final}}=\frac{1}{4}\left(Y_{1}+Y_{2}+Y_{3}+Y_{4}\right).$
This fused representation is then projected through a linear layer and passed to the decoder for segmentation. By enforcing directional consistency and resonance across scales, CR-MSM maintains the causal flow of information within the U-shaped architecture and produces multi-scale features that are semantically aligned, computationally efficient, and highly effective for medical image segmentation. The detailed procedure of the proposed CR-MSM is summarized in Algorithm 1. In which, $\{F_{i}\}_{i=1}^{N}$ are the encoder feature maps. Each $F_{i}$ is transformed into four directional views $\{X_{j}\}_{j=1}^{4}$, which are processed by SSM to produce $\{Y_{j}\}_{j=1}^{4}$. The fused output at scale $i$ is $\widetilde{F}_{i}$ through causal-resonance averaging, and $\{Z_{i}\}_{i=1}^{N}$ are the linearly projected skip features for the decoder.
Algorithm 1 Causal-Resonance Multi-Scale Module (CR-MSM)
1:Encoder feature maps $\{F_{i}\}_{i=1}^{N}$
2:Fused skip features $\{Z_{i}\}_{i=1}^{N}$
3:for $i\leftarrow 1$ to NN do
4:  Generate four views: $F_{i}^{\text{orig}},F_{i}^{\text{T}},F_{i}^{\text{flip}},F_{i}^{\text{flipT}}$
5:  Flatten each view to sequences: $X_{1},\ldots,X_{4}$
6:  Apply SSM: $Y_{j}\leftarrow\mathrm{SSM}(X_{j})$ for $j=1..4$
7:  Fuse by causal-resonance averaging:
$\widetilde{F}_{i}=\tfrac{1}{4}(Y_{1}+Y_{2}+Y_{3}+Y_{4})$
8:  Linear projection for skip connection: $Z_{i}\leftarrow\mathrm{Linear}(\widetilde{F}_{i})$
9:end for
10:return $\{Z_{i}\}_{i=1}^{N}$
@@ 4.1.2 Evaluation metrics
Here, $S_{P}$ denotes the set of surface points from the predicted segmentation, and $S_{G}$ denotes the set of surface points from the ground truth segmentation. The function $\text{dist}(x,S)$ represents the shortest distance from point $x$ to the surface $S$. The parameter $\tau$ is a predefined tolerance value, and $|\cdot|$ denotes the cardinality, i.e., the number of points in the set.
@@ 4.2 Quantitative and Qualitative results
On the BTCV dataset, our model outperforms the second-best model U-Mamba by 0.92% and 1.15% in DSC and NSD, respectively. We follow U-Mamba’s evaluation standard, eight out of thirteen organ classes (Aorta, Gallbladder, Kidney (L), Kidney (R), Liver, Pancreas, Spleen, and Stomach) are included in the calculation. Mamba based methods demonstrate strong overall performance, with our model maintaining a lead of 0.92% and 1.15% over U-Mamba in DSC and NSD. For the ACDC dataset, where all methods achieve relatively high accuracy, our method still surpasses SwinUNETR by 0.32% and 0.20% in DSC and NSD, respectively. On the EndoVis17 dataset, where both CNN-based and Mamba-based methods perform well, our method achieves further improvements, surpassing U-Mamba by 1.69% in DSC and 1.84% in NSD. Finally, on the ATLAS23 dataset, our model secures improvement by 1.27% and 1.68% in DSC and NSD, respectively while compare with U-Mamba.